OpenStudy (anonymous):

help in the su nlight, tick 7 ft tall casts a shadow 3ft long. At the same time a nearby tree casts a shadow 11 ft long. how tall is the three? round to the nearest tenth.

Directrix (directrix):

Triangle I: tan x = 7/3 tan x = 2.333 x = 66.8 degrees Note: The tangent function of an acute angle of a right triangle is defined to be the opposite leg length over the adjacent leg length ------- Triangle II: tan 66.8 degrees = T/11 where T is the length of the tree T = 11 (tan 66.8) T = 11 ( 2.333) T = 25.7 feet to the nearest tenth